# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ B.1 Proof of Lemma 1
Lemma B.1.
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
$P(x_{1},...,x_{k})=0$ (B.1)
as $\mathcal{X}\equiv\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$. Then, the set $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$, unless Equation (B.1) is identically true.
@@
If $k=1$, the proof is trivial. Suppose $k>1$. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},...,x_{k,0})\in\mathbb{R}^{k}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, the set $\mathcal{X}$ is an $k-1$-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, it follows that the set $\mathcal{X}$ is an $k-1$-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{k}$ in the region where the gradient of $P(x_{1},...,x_{k})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P(x_{1},...,x_{k})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{k}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$. ∎
@@
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
@@
Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$ for each $t$. Hence, the union $\cup_{t=1}^{\infty}S_{t}$ is a union of a countable number of measure zero sets and thus itself has a Lebesgue measure zero. It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
B.2 Proof of Lemma 2
@@
$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$, $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.2), we obtain
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
@@
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0,$ (B.4)
@@
as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (B.3), we get
$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.5)
@@
Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most one column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$), so $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
@@
Now, from the invertibility of $C$ it follows that each column of $C$ as exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element. Therefore, there exists a permutation matrix $P$ and a diagonal matrix $D=\text{diag}(d_{1},...,d_{d})$ with nonzero diagonal elements such that $C=DP$. Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.

### v5

@@ C.1 Proof of Lemma 1
Lemma C.1.
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, i=1,,ki=1,...,k, that is not identically zero. Then, the set $\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$ has Lebesgue measure zero in $\mathbb{R}^{k}$.
@@
Proof by induction on $k$. For $k=1$, the claim clearly holds, as a nontrivial univariate polynomial has finitely many zeros. For the induction step, suppose the set $\{(x_{1},...,x_{k-1})\in\mathbb{R}^{k-1}|P(x_{1},...,x_{k-1})=0\}$ has Lebesgue measure zero in $\mathbb{R}^{k-1}$. Write a point $x\in\mathbb{R}^{k}$ as $x=(x^{*},z)$ where $x^{*}=(x_{1},...,x_{k-1})\in\mathbb{R}^{k-1}$ and $z\in\mathbb{R}$ is the last component. Then, we can view $P(x^{*},z)$ as a polynomial in $z$, whose coefficients are polynomials in $x^{*}$:
$P(x^{*},z)=\sum_{i=0}^{s}a_{i}(x^{*})z^{i},$ (C.1)
where $s$ is the degree of $P$ in the variable $z$ and each coefficient $a_{i}(x^{*})$ is itself a polynomial in k1k-1 variables.
Define $\mathcal{E}\equiv\{x^{*}\in\mathbb{R}^{k-1}:a_{0}(x^{*})=a_{1}(x^{*})=\cdots=a_{s}(x^{*})=0\}$. If $x^{*}\in\mathcal{E}$, then all coefficients vanish, so $P(x^{*},z)=0$ for all $z$. If $x^{*}\notin\mathcal{E}$, then $P(x^{*},z)$ is a nontrivial univariate polynomial in $z$. Because $P$ is not identically zero, at least one coefficient $a_{i}(x^{*})$, $i\in\{1,...,s\}$, is nontrivial. Then, by the induction hypothesis, the zero set of this nontrivial $a_{i}(x^{*})$ (and hence $\mathcal{E}\subset\{x^{*}:a_{i}(x^{*})=0\}$) has measure zero in $\mathbb{R}^{k-1}$. Also, for each fixed $x^{*}\in\mathbb{R}^{k-1}$, define the set $\mathcal{S}(x^{*})\equiv\{z\in\mathbb{R}:P(x^{*},z)=0\}$. If $x^{*}\notin\mathcal{E}$, then $P(x^{*},z)$ is a nontrivial univariate polynomial, so $\mathcal{S}(x^{*})$ is a finite set of points in $\mathbb{R}$ and thereby has measure zero. If $x^{*}\in\mathcal{E}$, then $\mathcal{S}(x^{*})=\mathbb{R}$.
Now, the zero set of $P$ is $\mathcal{Z}\equiv\{(x^{*},z)\in\mathbb{R}^{k}:P(x^{*},z)=0\}$. Denoting by $\lambda_{k}(\cdot)$ the $k$-dimensional Lebesgue measure, we have by Fubini’s theorem:
$\lambda_{k}(\mathcal{Z})=\int_{\mathbb{R}^{k-1}\times\mathbb{R}}\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}dx^{*}dz=\int_{\mathbb{R}^{k-1}}\left(\int_{\mathbb{R}}\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}dz\right)dx^{*},$ (C.2)
where $\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}$ is an indicator function that takes the value one if $(x^{*},z)\in\mathcal{Z}$ and the value zero otherwise. The inner integral on the right side of (C.2) is the one-dimensional Lebesgue measure of the slice $\{z:(x^{*},z)\in\mathcal{Z}\}=S(x^{*})$, so $\lambda_{k}(\mathcal{Z})=\int_{\mathbb{R}^{k-1}}\lambda_{1}(S(x^{*}))dx^{*}$. If $x^{*}\notin\mathcal{E}$, the slice $S(x^{*})$ is finite, implying that $\lambda_{1}(S(x^{*}))=0$. On the other hand, if $x^{*}\in\mathcal{E}$, the slice equals $\mathbb{R}$, implying that $\lambda_{1}(S(x^{*}))=\infty$. However, since $\mathcal{E}$ has measure zero in $\mathbb{R}^{k-1}$, it follows that $\lambda_{1}(S(x^{*}))$ is zero almost everywhere, and hence, $\lambda_{k}(\mathcal{Z})=0$, concluding the proof. ∎
@@
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that for all $\alpha_{1,t},...,\alpha_{M,t}$ there exists (invertible) matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible, e.g., $B_{1}=...=B_{M}=I_{d}$, so $B_{y,t}=I_{d}$ (as $\sum_{m=1}^{M}\alpha_{m,t}=1$ for all $t$). The set of points $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ for which $B_{y,t}$ is singular is $S_{t}\equiv\{([B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma C.1 that $S_{t}$ has Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
@@
Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has Lebesgue measure zero in $\mathbb{R}^{d\times dM}$ for each $t$. Hence, the union $\cup_{t\in\mathbb{N}}S_{t}$ is a union of a countable number of measure zero sets and thus itself has Lebesgue measure zero. It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
C.2 Proof of Lemma 2
@@
$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (C.3)
where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$; $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$; $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$, i=1,,pi=1,...,p, are the time-varying AR matrices; and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (C.3) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, m=1,,Mm=1,...,M, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (C.3), we obtain
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (C.4)
@@
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0$ (C.5)
@@
almost surely, as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (C.4), we get
$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (C.6)
@@
Since $B_{y,t}$ is invertible almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most one column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$), so $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
@@
Now, from the invertibility of $C$ it follows that each column of $C$ has exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element. Therefore, there exists a permutation matrix $P$ and a diagonal matrix $D=\text{diag}(d_{1},...,d_{d})$ with nonzero diagonal elements such that $C=DP$. Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.